OpenStudy (anonymous):

y-3x=0 need to solve to gets points and graph

OpenStudy (anonymous):

Just plug in any arbitrary number for one of the variables, and see what the other variable's value is. y - 3x = 0 x = 0 y - 0 = 0 y = 0 (0,0) y= 6 6 - 3x = 0 x = 2 (2,6)

OpenStudy (anonymous):

y - 3x = 0 y = 3x plug in numbers to make a graph y = 3(1) = 3 y = 3(2) = 6 now go back and find y 3 = 3x = 1 6 = 3x = 2 so two points on the graph are (1,3) and (2,6)